OpenStudy (anonymous):

How many 5/8s are in 3?

OpenStudy (anonymous):

3/(5/8) =3*8/5 =24/5 =4.8 hence there are four 5/8s in 3

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

(Y)